OpenStudy (anonymous):

solve the equation please help 6(x-0.8)-0.2(5x-4)=6 Options A=-0.5 B=-2 C=0.5 D=2 Please show work

OpenStudy (anonymous):

use distributive property: \[6(x-0.8)-0.2(5x-4)=6\]\[6x-4.8-x+0.8=6\]combine like terms:\[5x-4=6\]add four two both sides:\[5x=10\]divide both sides by five:\[x=2\] \[Answer- D:~ 2\]

OpenStudy (anonymous):

@tabbycat737 helpful? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but wat i tried to do is the answers ther i put them into the problem to see if it will =6 is that good too

OpenStudy (anonymous):

its easier to just work through one problem rather than guess and check but it really depends on the situation...you should just make sure that if you're asked to show work, your not caught off guard :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anytime! :)